Let the market supply function of 2000 sellers is given as Qs = 4000 + 20Px. and the individual demand function of 5000 buyers is given as 2px = 1500 – 0.2Qd. Then compute
The market clearing price.
The equilibrium quantity.
Explain graphically.
1
Expert's answer
2021-10-05T09:21:37-0400

Solution:

At equilibrium: Qd = Qs

Qd = 7500 – 10px

7500 – 10px = 4000 + 20px

7500 – 4000 = 20px + 10px

3500 = 30px

Px = 117

 

The market-clearing price = 117

 

Substitute to derive the equilibrium quantity:

Qd = 7500 – 10(117) = 7500 – 1170 = 6330 units

 

This is graphically displayed as follows: